OpenStudy (anonymous):

3log(subscript6)(x+1)=9

OpenStudy (anonymous):

what do i do with the 3

ganeshie8 (ganeshie8):

\[\large 3\log_6 (x+1) = 9\]

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

thats a good question :) get rid off it.

ganeshie8 (ganeshie8):

divide 3 both sides

OpenStudy (anonymous):

6^9=(x+1)^3?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

taht just made my day easier lol

OpenStudy (anonymous):

nope wait

OpenStudy (anonymous):

was i right?

OpenStudy (anonymous):

3 is basically multiplied with the log so \[\log_{6}(x+1)=9/3\] \[\log_{6}(x+1)=3\]

OpenStudy (anonymous):

Now you can do it the way i told u earlier

OpenStudy (anonymous):

k

OpenStudy (anonymous):

ty